OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 3) (3, 7)

OpenStudy (amistre64):

how do we determine the slope between 2 points?

OpenStudy (amistre64):

slope is measured by how far our y changes with respect to changes in our x value

OpenStudy (anonymous):

m = (y2 -y1)/(x2 - x1)

OpenStudy (amistre64):

the x value moves from 7 to 3; thats a change of -4 the y value moves from 3 to 7; thats a change of +4 our ratio then is: change in y / change in x = 4/-4, simplified to: -1

OpenStudy (amistre64):

y-Py = m(x-Px) y = mx -mPx +Py ; where P(x,y), and m=slope

OpenStudy (anonymous):

wow, that equation takes me back a couple years

OpenStudy (amistre64):

y = -x +1(3)+7